OpenStudy (anonymous):

\huge lim_{x rightarrow 1} frac{sqrt[5]{2x+1}-1}{sqrt[4]{2x+1}-1}

myininaya (myininaya):

\[\lim_{x \rightarrow 1}\frac{\sqrt[5]{2x+1}-1}{\sqrt[4]{2x+1}-1}\] Correct?

myininaya (myininaya):

Is so the function is continuous at x=1 so just plug it right on in :)

myininaya (myininaya):

if not is *

OpenStudy (anonymous):

@myininaya : sry i wrote it wrong ,Limit Tends to 0 .

myininaya (myininaya):

Ok and hey which way do you prefer to do this? algebraically or l'hospital?

OpenStudy (anonymous):

algebraically

myininaya (myininaya):

\[\lim_{x \rightarrow 0}\frac{\sqrt[5]{2x+1}-1}{\sqrt[4]{2x+1}-1}\] Let me think on the algebraic approach I think it might be a substitution So the lcm of 5 and 4 is 20 So lets try the sub \[u^{20}=2x+1\] So that means \[(u^{20})^\frac{1}{5}=(2x+1)^\frac{1}{5} => u^4=\sqrt[5]{2x+1}\] and \[(u^{20})^\frac{1}{4}=(2x+1)^\frac{1}{4} => u^5=\sqrt[4]{2x+1}\]

myininaya (myininaya):

So if x goes to 0 then what does u go to ?

myininaya (myininaya):

\[u^{20}=2x+1 => u=(2x+1)^\frac{1}{20} \text{ correct?}\] So can you tell me what u goes to if x goes to 0?

myininaya (myininaya):

Eyad we are almost there :)